OpenStudy (anonymous):

a car produces propulsion force 12 kilo N. when it is starling at 25m per second, what is the power developed by the car's engine?

OpenStudy (michele_laino):

the requested power P, is given by the subsequent formula: \[\Large P = {\mathbf{F}} \cdot {\mathbf{v}} = Fv = 12 \times {10^3} \times 25 = ...watt\]

OpenStudy (shamim):

We know Power p=w/t

OpenStudy (shamim):

w=Fs

OpenStudy (shamim):

p=Fs/t=F*s/t=Fv